OpenStudy (anonymous):

Find all critical points of R(t)=3t+77-15ln(t^2-3)

OpenStudy (amistre64):

R' = 3 -(30t)/(t^2-3) i think

OpenStudy (amistre64):

3t^2 -30t -9 ------------- t^2- 3

OpenStudy (amistre64):

3(t^2 -10 -3) would be the top if I did it right....

OpenStudy (amistre64):

-10t....

OpenStudy (anonymous):

I Got R'=3-15/(t^2-3)(2t)

OpenStudy (amistre64):

yep, as long as thats just screen clutter and not a glaring error :)

OpenStudy (amistre64):

add the parts like fractions, 3(t^2 -3) -30t ------------ t^2 -3

OpenStudy (anonymous):

Sorry how did you get 3(t^2-3) on top?

OpenStudy (amistre64):

the top becomes a quadratic.... a fancy name for "highest power is a 2"

OpenStudy (amistre64):

ok... you pretty much have the form: 30 3 - ---- now how do we add fractions togther? b

OpenStudy (anonymous):

You have to have a common denominator?

OpenStudy (anonymous):

But if R'(t)=3- 15 (2t) ----- (t^2 -3)

OpenStudy (anonymous):

That would be 3 - 30t ------- (t^2 - 3)

OpenStudy (amistre64):

30t 3 - ------- is what we got right? (t^2 -3)

OpenStudy (anonymous):

That's what I got

OpenStudy (amistre64):

then to add the "fractions" mulitply the left part by: 3 t^2-3 -- x ----- right? 1 t^2-3

OpenStudy (anonymous):

Oh! Okay

OpenStudy (amistre64):

:) deriving...no problem.... fractions? ehhhh :)

OpenStudy (amistre64):

3t^2 -30t -9 ----------- t^2 -3

OpenStudy (amistre64):

quadratic formula for the top to find the critical points...

OpenStudy (amistre64):

5 +- [sqrt(900 - (4)(3)(-9))]/6

OpenStudy (anonymous):

Where is the 5 from?

OpenStudy (amistre64):

5 +- [sqrt(1008)]/6

OpenStudy (amistre64):

30/6 -b/2a

OpenStudy (amistre64):

you familiar with the quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

good heres what I get: t = 5 +- 2sqrt(63)/3

OpenStudy (anonymous):

Did you do quadratic formula of 3t^2-30t-9?

OpenStudy (amistre64):

yep..... as long as I kept track of myself

OpenStudy (amistre64):

30 sqrt(900-(4)(3)(-9)) --- +- -------------- 6 6

OpenStudy (anonymous):

So 30 +- sqrt(900+108) ---------------------- 6

OpenStudy (amistre64):

we can simplify that by factoring out a 3 first.... sneaky little bugger it is :)

OpenStudy (amistre64):

3(t^2 -10t -3) now quad the left side

OpenStudy (amistre64):

10 +- sqrt(100 -(4)(-3)(1)) ----------------------- 2

OpenStudy (amistre64):

10 +- sqrt(112) --- -------- same result, just easier to see 2 2

OpenStudy (amistre64):

112 = 16*7 4sqrt(7)

OpenStudy (anonymous):

So 5 +- 2sqrt(7)

OpenStudy (amistre64):

thats better :)

OpenStudy (amistre64):

you know how to test for min and max?

OpenStudy (anonymous):

So those are the critical points? The neg and pos value from that?

OpenStudy (anonymous):

No

OpenStudy (amistre64):

yep... those are the critical points. to see how they are behaving, you take a second derivative....derive the derivative.

OpenStudy (anonymous):

I don't think I have to do that.

OpenStudy (amistre64):

you sure? its fun :) if you have to determine if they are MIN MAX or inflections, then we would, but if you just gotta find criticals, thats it.... dont forget to include the end points of the graph as well, they might be important

OpenStudy (anonymous):

Wow, thanks! Can you help me with one more?

OpenStudy (amistre64):

i dunno....... i feel a bout of stupidy brewing amongst me brain cells :) ........sure why not :)

OpenStudy (anonymous):

Find all critical points of P(t)=-cos(4t)+39-2t

OpenStudy (anonymous):

Holy cow sorry my computer spazzed out

OpenStudy (amistre64):

thats just the cosine function moved about.... should be rather simple to determine

OpenStudy (anonymous):

I have found P'(t) if that helps

OpenStudy (amistre64):

P' = 4sin(4t) - 2 right?

OpenStudy (anonymous):

Yes that's what I got

OpenStudy (amistre64):

good, then we are either both genuises or idiots :) Ill go with genuisesss

OpenStudy (anonymous):

(=

OpenStudy (amistre64):

0 = 4sins(4t) -2 solve for sin(4t)

OpenStudy (anonymous):

sin(4t)=1/2

OpenStudy (amistre64):

what does 4t have to equal for sin to be 1/2?

OpenStudy (amistre64):

there are 2 angles, that will satisfy this

OpenStudy (anonymous):

I have no idea.

OpenStudy (amistre64):

what? you gotta have SOME idea.... its just a circle :)

OpenStudy (amistre64):

sin(30) = 1/2 and sin(150) = 1/2

OpenStudy (anonymous):

How do I figure the two angles that will satisfy that?

OpenStudy (amistre64):

you can either remember back to trig that a 30-60-90 triangle has sides of 1-2-sqrt(3) and interpolate from that, OR use the sin inverse function on a calculator

OpenStudy (anonymous):

so sin^-1(1/2)=4t

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

what from that tells me that 30 and 150 are what I'm looking for?

OpenStudy (anonymous):

Sorry I want to understand

OpenStudy (amistre64):

sin^-1 will only give you one angle tho...30. You have to remember to cross over the yaxis to get to the other one...150

OpenStudy (amistre64):

4t = 30 and 4t = 150 solve for t :)

OpenStudy (amistre64):

might be better to use radians instead of degrees, half dozen of one or 6 of the other

OpenStudy (anonymous):

Oh! Okay thanks. Moving on

OpenStudy (amistre64):

4t = pi/6 4t = 5pi/6

OpenStudy (anonymous):

whoa where did that come from?

OpenStudy (amistre64):

what!!! you see a bee??

OpenStudy (anonymous):

No I mean the 4t=pi/6 and 4t=5pi/6

OpenStudy (anonymous):

Wait is that 30 degrees and 150 degrees in radians?

OpenStudy (amistre64):

:) there are a few ways to express angles. you can use degrees: 30 and 150 are degrees. OR you can use "radians" which measures the angle using the radius itself as a measuring stick. a circle is 360 degrees around OR 2pi radians pi/6 = 30 degrees 5pi/6 = 150 degrees

OpenStudy (anonymous):

Okay

OpenStudy (amistre64):

To get a number for "t" we should use radians :)

OpenStudy (anonymous):

Okay

OpenStudy (amistre64):

t = pi/24 t = 5pi/24 is what I come up with

OpenStudy (amistre64):

30/4 or 150/4 depending on what your answers are looking for

OpenStudy (anonymous):

so t=.1309 + pin/2

OpenStudy (amistre64):

t = pi/24 = .1309 or t = 5pi/24 = .6545

OpenStudy (amistre64):

for degrees... t = 7.5 degrees or t = 37.5 degrees

OpenStudy (anonymous):

Okay. I have a question.

OpenStudy (amistre64):

go ahead and ask :)

OpenStudy (anonymous):

usually before solving for t my professor asks "which quadrants is sin negative in" why did we not have to look for that in this problem?

OpenStudy (amistre64):

sin(angle) = 1/2 sin is only positive in Q1 and Q2 .... no need to look any where else, unless you can think of a reason to..

OpenStudy (anonymous):

Okay. That makes sense. Thank you sooo much!

OpenStudy (amistre64):

when sin = -1/2, it puts a different value into the equation that is not a critical point

OpenStudy (amistre64):

you are soooo welcome :)

OpenStudy (amistre64):

if you have any more questions, just start a new posting :)